Factor the trinomial completely. If this trinomial contains a greatest common factor (other than 1), dont forget to factor out the GCF first x +xy -56y Select the correct choice below and, if necessary, fill in the answer box to complete your choice OA, x2 + xy-56y2 = ! O B. x2 + xy-56y2 is prime 2
Answer #1

x^2+xy-56y^2

=\left(x^2-7xy\right)+\left(8xy-56y^2\right)

r (x - 7y) + 8y (x - 7y)

=(x-7y)(x+8y)